# DeepScholar-Bench: A Live Benchmark and Automated Evaluation for Generative Research Synthesis

### v1

@@ 8.2.1. DeepResearcher
For our experiments, we used Llama-4-Scout-17B-16E-Instruct as the base model, replacing the originally proposed DeepResearcher-7b, since it demonstrated consistently better retrieval-augmented reasoning performance in our experiments. The prompt was slightly modified to align with LLama-4 prompt style as detailed in Box 8.2.1. The retrieval depth was set to 10 sources per query, which is the default in the system and provides a balanced trade-off between coverage and efficiency. We restricted each query to a single rollout with a maximum of 10 steps, following the DeepResearcher defaults; this limit is generous as most rollouts converge in fewer than three steps, but it ensures the system has headroom for more complex queries. The default web search API was replaced with ArXiv search to comply with our benchmark settings.
8.2.2. Openscholar
@@
For consistency with other baselines, we employ the Llama-4-Scout-17B-16E-Instruct model for generation. The retrieval pipeline initially collects 100 text segments from peS2o_v3 using the default pes2o_contriever model. The reranker used is OpenScholar_Reranker, also kept at its default setting. To align parameterization across baselines, we increase the number of sources used in generation (top_n) from 10 to 30. Furthermore, the default search API is replaced with the arXiv API, to provided a controlled retreival corpus and API in our experiments.
8.2.3. Search AI
Search AIs are implemented using the open-source OpenDeepSearch (ODS) framework that enables deep web search and retrieval. In particular, the ODS ReAct Agent (instantiated from the smolagents.ToolCallingAgent) is used along with the search agent as an external tool. At each reasoning step, the ReAct agent can either invoke the search agent through a web_search action or decide to produce a final_answer. The search agent interfaces with the search API to fetch relevant academic articles given a query, after which an LLM generates concise summaries of the retrieved content. To maintain consistency with the benchmark setting, the standard search API was replaced with the arXiv API. The regular search agent fails when tasked with full abstract queries; hence the ReAct-based agent was employed, which generates shorter, more effective searchable queries. The agent keeps track of retrieved results across turns, allowing references to past evidence during the reasoning process. After a maximum of 5 iterations, the agent is compelled to conclude with a final response, ensuring bounded computational steps.
For the parameterization of the Search AIs, we set the search agent to retrieve 30 results per query, which is more generous than the default in order to establish fair comparability with other baselines. The reranker parameter was left at infinity, aligning with its default configuration, to avoid prematurely constraining candidate results. The maximum iteration limit was fixed at 5, consistent with the default setup of the ODS framework, providing sufficient exploration without excessive search depth. The ReAct prompt was slightly modified to tailor to the specific use of the ArXiv search API, as presented in Box 8.2.3, 8.2.3 and 8.2.3.
8.2.4. STORM
@@
8.2.5. OpenAI’s DeepResearch
We use OpenAI’s DeepResearch system based on the o3-deep-research model with a custom MCP to only search ArXiv and return results per query. To prevent the model from getting search results after the given paper was uploaded, the MCP used a custom endpoint to set the latest date that it should retrieve. All other settings were set to default values.
8.3. DeepScholar-Base Details and Configurations
@@
- Retrieval:
In this stage, an LLM generates search queries conditioned on the input abstract and summaries of prior retrievals. Each query is submitted to the configured search API (ArXiv, tavily, etc.) to obtain up to relevant papers within the specified date range. The code and prompt used for this step are provided in Figure 5 and Box 8.3 respectively. This process is repeated times.
@@
Figure 5. Retrieval stage: query generation and batched search. - Filtering:
Retrieved results are refined using two semantic operators from LOTUS (Patel et al., 2025; noa, 2025b): Sem-Filter and Sem-TopK, which together select the top most relevant papers. The code is given in Figure 6.
@@
Figure 6. Sem-Filter and Sem-TopK code for Filtering Step in DeepScholar-base - Final Generation:
The filtered set of papers is then aggregated via a Sem-Agg query to produce the final output. The corresponding code for this step is shown in Figure 7 with prompt in Box 8.3.
@@
Figure 7. Sem-Agg for final generation in DeepScholar-base
@@
8.4. Evaluation Details
8.4.1. Reference Coverage
8.4.2. Ablation Study on Verifiability
8.4.3. Document Importance Across Human Exemplars
In this section, we illustrate the distribution of document importance, measured by the number of citations of references in the human-written exemplars in DeepScholar-Bench. Figure 11 reports two histograms: (a) the distribution of citation counts across all references, and (b) the distribution restricted to references that appear on ArXiv. We plot the logarithm of citation counts, with values obtained from the OpenAlex API (noa, [n.d.]j), an open and widely used scholarly database that provides citation-level metadata. While citation counts in OpenAlex may not exactly match those from other sources such as Google Scholar, the relative counts are consistent, making it a reliable open-source alternative.

### v2

@@ A.5.1 DeepResearcher
For our experiments, we used Llama-4-Scout-17B-16E-Instruct as the base model, replacing the originally proposed DeepResearcher-7b, since it demonstrated consistently better retrieval-augmented reasoning performance in our experiments. The prompt was slightly modified to align with LLama-4 prompt style as detailed in Box A.5.1. The retrieval depth was set to 10 sources per query, which is the default in the system and provides a balanced trade-off between coverage and efficiency. We restricted each query to a single rollout with a maximum of 10 steps, following the DeepResearcher defaults; this limit is generous as most rollouts converge in fewer than three steps, but it ensures the system has headroom for more complex queries. The default web search API was replaced with ArXiv search to comply with our benchmark settings.
A.5.2 Openscholar
@@
For consistency with other baselines, we employ the Llama-4-Scout-17B-16E-Instruct model for generation. The retrieval pipeline initially collects 100 text segments from peS2o_v3 using the default pes2o_contriever111https://huggingface.co/akariasai/pes2o_contriever model. The reranker used is OpenScholar_Reranker222https://huggingface.co/OpenSciLM/OpenScholar_Reranker, also kept at its default setting. To align parameterization across baselines, we increase the number of sources used in generation (top_n) from 10 to 30. Furthermore, the default search API is replaced with the arXiv API, to provide a controlled retrieval corpus and API in our experiments.
A.5.3 Search Agent
Search Agents are implemented using each LLM API with search API access. We implement a ReAct Agent, instantiated from the smolagents.ToolCallingAgent, with a system prompt based on the open-source OpenDeepSearch (ODS) framework designed for deep web search and retrieval, along with the search agent as an external tool. At each reasoning step, the ReAct agent can either invoke the search agent through a web_search action or decide to produce a final_answer. The search agent interfaces with the search API to fetch relevant academic articles given a query, after which an LLM generates concise summaries of the retrieved content. To maintain consistency with the benchmark setting, the standard search API was replaced with the arXiv API. The regular search agent fails when tasked with full abstract queries; hence the ReAct-based agent was employed, which generates shorter, more effective searchable queries. The agent keeps track of retrieved results across turns, allowing references to past evidence during the reasoning process. After a maximum of 5 iterations, the agent is compelled to conclude with a final response, ensuring bounded computational steps.
For the parameterization of the Search Agents, we set the search agent to retrieve 30 results per query, which is more generous than the default in order to establish fair comparability with other baselines. The maximum iteration limit was fixed at 5, consistent with the default setup of the ODS framework, providing sufficient exploration without excessive search depth. The ReAct prompt was slightly modified to tailor to the specific use of the ArXiv search API, as presented in Box A.5.3, A.5.3 and A.5.3.
A.5.4 STORM
@@
A.5.5 OpenAI’s DeepResearch
We use OpenAI’s DeepResearch system based on the o3-deep-research333https://platform.openai.com/docs/models/o3-deep-research model with a custom MCP to only search ArXiv and return results per query. To prevent the model from getting search results after the given paper was uploaded, the MCP used a custom endpoint to set the latest date that it should retrieve. All other settings were set to default values.
A.6 DeepScholar-ref Details and Configurations
@@
- Retrieval
In this stage, an LLM generates search queries conditioned on the input abstract and summaries of prior retrievals. Each query is submitted to the configured search API (ArXiv, tavily, etc.) to obtain up to relevant papers within the specified date range. The code and prompt used for this step are provided in Figure 8 and Box A.7 respectively. This process is repeated times.
@@
Figure 8: Retrieval stage: query generation and batched search. - Filtering
Retrieved results are refined using two semantic operators from LOTUS [37, 28]: Sem-Filter and Sem-TopK, which together select the top most relevant papers. The code is given in Figure 9.
@@
Figure 9: Sem-Filter and Sem-TopK code for Filtering Step in DeepScholar-ref - Final Generation
The filtered set of papers is then aggregated via a Sem-Agg query to produce the final output. The corresponding code for this step is shown in Figure 10 with prompt in Box A.7.
@@
Figure 10: Sem-Agg for final generation in DeepScholar-ref
@@
A.7 Examples of Generated Reports
We provide examples of generated reports from different systems for paper 0 in our dataset (according to Table 5 i.e., ‘TaxAgent: How Large Language Model Designs Fiscal Policy’ with Arxiv ID of 2506.02838. Figures LABEL:fig:example:ours, 11, 12, 13 and 14 display the generated reports produced by DeepScholar-ref, Search Agent, DeepResearcher, OpenScholar, and Storm respectively, all using Llama-4.